A cashier has P 50, P 100, and P500 bills. the number of P50 bills is 5 more than 7 times the number of P100 and the number of P500 bills is four-fifths the number of P100 bills. If the total amounts that she has P18 100, how many of each type of bills are there?
please give me your representation...
thank you

Answer by stanbon(75887)

cashier has P 50, P 100, and P500 bills. the number of P50 bills is 5 more than 7 times the number of P100 and the number of P500 bills is four-fifths the number of P100 bills. If the total amounts that she has P18 100, how many of each type of bills are there?
--------------
Let # of P50 be "x":
Let # of P100 be "y":
Let # of P500 be "z":
--------------------------
EQUATIONS:
x = 7y + 5
z = (4/5)y
50x + 100y + 500z = 18100
-------------------------------
Rearrange:
50x + 100y + 500z = 18100
x -7y + 0 = 5
0 + 4y - 5z = 0
----------------------------
Use any method you know to get:
x = 152
y = 21
z = 16
==================
Cheers,
Stan H.